OpenStudy (anonymous):

Verify the identity. sin 4u = 2 sin 2u cos 2u

OpenStudy (anonymous):

A rule of thumb for something like this is to always start on the side that seems the most complicated

OpenStudy (anonymous):

2cos(x) sin(x) = 2sin(2x) use that

OpenStudy (anonymous):

I'm really bad at these..

OpenStudy (anonymous):

Use sourwing 's equation to help you since that is a trig identity that really helps

OpenStudy (anonymous):

I'm not sure what to do next

OpenStudy (anonymous):

..sin(4u) = sin(2u+2u) = sin(2u)cos(2u) + cos(2u)sin(2u) = 2 sin(2u) cos(2u). ?

OpenStudy (anonymous):

2 cos(2u) sin(2u) = 2 sin(2 * 2u) = 2 sin(4u)

OpenStudy (anonymous):

yes, you can do it that way

OpenStudy (anonymous):